# Harmonia: A Multi-Agent Reinforcement Learning Approach to Data Placement and Migration in Hybrid Storage Systems

### v1

@@ Abstract.
Hybrid storage systems (HSS) combine multiple storage devices with diverse characteristics to achieve high performance and capacity at low cost. The performance of an HSS highly depends on the effectiveness of two key policies: (1) the data-placement policy, which determines the best-fit storage device for incoming data, and (2) the data-migration policy, which rearranges stored data across the devices to sustain high HSS performance. Prior works focus on improving only data placement or only data migration in HSS, which leads to sub-optimal HSS performance. Unfortunately, no prior work tries to optimize both policies together.
Our goal is to design a holistic data-management technique for HSS that optimizes both data-placement and data-migration policies to fully exploit the potential of an HSS. We propose Harmonia, a multi-agent reinforcement learning (RL)-based data-management technique that employs two light-weight autonomous RL agents, a data-placement agent and a data-migration agent, which adapt their policies for the current workload and HSS configuration, and coordinate with each other to improve overall HSS performance.
@@ 1. Introduction
A hybrid storage system (HSS) (xiao2016hs, ; wang2017larger, ; lee2014mining, ; felter2011reliability, ; bu2012optimization, ; krish2016efficient, ; lin2011hot, ; niu2018hybrid, ; tai2015sla, ; ou2014edm, ; cheng2015amc, ; zong2014faststor, ; matsui2017design, ; heuristics_hyrbid_hystor_sc_2011, ; lv2013probabilistic, ; guerra2011cost, ; elnably2012efficient, ; heuristics_usenix_2014, ; zhang2010automated, ; wu2012data, ; iliadis2015exaplan, ; lv2013hotness, ; matsui2017tri, ; feng2014hdstore, ; yang2017autotiering, ; kim2011hybridstore, ; liu2019hierarchical, ; luo2020optimal, ; singh2022sibyl, ; doudali2019kleio, ; ren2019archivist, ; cheng2019optimizing, ; shetti2019machine, ) is a cost-effective solution for both high-performance and large-capacity storage requirements of data-intensive applications such as databases (chan-signmod-1998, ; FastBitA9, ; redis-bitmaps, ), graph processing (beamer-SC-2012, ), and genome analysis (lander-nature-2001, ; altschul-jmb-1990, ; myers-jacm-1999, ). HSS combine multiple storage devices (e.g., solid-state drives (SSDs) (inteloptane, ; intels4510, ; samsung2017znand, ; samsung-980pro, ; samsungmlc, ), hard disk drives (HDDs)) of diverse characteristics (e.g., I/O bandwidth, storage capacity). A typical HSS consists of fast storage devices with limited capacities and slow storage devices with large capacities.
@@
The performance of an HSS is highly dependent on two key policies, data placement and data migration. The data-placement policy improves HSS performance by determining the best-fit storage device in the HSS for incoming I/O requests. The data-migration policy rearranges data across the storage devices to sustain high HSS performance, which can degrade over time due to (1) misplacement of data, and (2) changes in workload access patterns.
@@
We identify four key challenges in designing efficient data-place-ment and data-migration policies. First, in data-intensive environments, workload access patterns and HSS conditions (e.g., access latencies, device capacity utilization) can change frequently, which makes it hard to optimize the placement and migration policies. Second, the data-placement policy should have a low performance overhead because it operates in the critical path of I/O request handling. Third, the data-migration policy needs to migrate data across the storage devices in a timely manner without impacting the latency of incoming I/O requests. Fourth, the data-placement and data-migration policies should not make conflicting decisions, which can adversely impact HSS performance and device lifetimes.
@@
Limitations of prior works. A large body of prior work proposes various techniques to optimize either data-placement (e.g., (singh2022sibyl, ; matsui2017design, ; matsui2017tri, ; ren2019archivist, ; heuristics_hyrbid_hystor_sc_2011, ; wu2012data, ; wu2019data, ; yang2014statistics, ; yolchuyev2021data, ; strunk2012hybrid, )) or data-migration (e.g., (shetti2019machine, ; doudali2019kleio, ; lin2011hot, ; lin2017efficient, ; lin2018buffer, )) policies, but these prior works provide sub-optimal HSS performance. Our motivational study on performance-optimized (cost-optimized) HSS configurations demonstrates that even the state-of-the-art data-placement technique for HSS, Sibyl (singh2022sibyl, ), merely achieves 49.9% (41.4%) of the I/O performance of an Oracle policy that makes data-placement and data-migration decisions based on perfect knowledge of future access patterns (see §3 for more details). Based on our evaluation of various prior techniques, we identify two key reasons for such large performance gaps from the Oracle policy. First, no prior work proposes a holistic approach to optimize both data-placement and data-migration policies together. Second, heuristic- and supervised-learning-based policies (e.g., (singh2022sibyl, ; matsui2017design, ; matsui2017tri, ; doudali2019kleio, ; shetti2019machine, ; ren2019archivist, ; lin2011hot, ; lin2017efficient, )) require hand-tuning to provide performance benefits and may not adapt to changes in workload characteristics and HSS states.
@@
Since there are no prior works that focus on optimizing both data-placement and data-migration policies, we combine prior data-placement and data-migration techniques to create four extended data-management techniques (see §3 for more details). Our evaluation shows that these extended techniques underperform because their data-placement and data-migration policies are taken from two different prior approaches and are oblivious to each other’s decisions. We conclude that for high HSS performance, both placement and migration policies need to be adaptive and coordinate with each other.
@@
Our goal is to develop a new, holistic data-management technique for HSS that optimizes both data-placement and data-migration policies in a coordinated way to improve the overall HSS performance. To this end, we propose Harmonia, [^1] Harmonia (wiki:harmonia, ) is the Greek goddess of harmony and balance, who brings peace to conflicts. a multi-agent online reinforcement learning (RL) based holistic approach for data placement and data migration in HSS.
Why RL? We choose RL for HSS data management because it offers three key benefits compared to heuristic- and supervised-learning-based approaches. First, RL dynamically adapts its decision-making to changes in workload access patterns and HSS conditions based on system-level feedback without human intervention. Second, RL can learn strategies that maximize long-term returns in complex tasks. Third, RL offers easy extensibility to various workloads and HSS configurations with minimum designer effort.
@@
Need for multi-agent RL. Despite the benefits of an RL-based approach, a single RL agent may not effectively (1) co-optimize (yang2020multi, ; yu2020meta, ; teh2017distral, ; sun2020adashare, ) different objectives like data placement and data migration, and (2) perform both tasks concurrently. Our motivational study demonstrates that a single-agent RL approach that performs both data placement and data migration shows 18.8% (27.7%) lower performance on average than Sibyl on a performance-optimized (cost-optimized) HSS (see §3 for more details).
@@
Key Idea. Harmonia is a multi-agent data-management approach in which two light-weight autonomous RL agents, a data-placement agent and a data-migration agent, improve their policies continuously, and coordinate with each other to improve HSS performance. First, we design Harmonia’s data-migration agent to identify migration candidates and their target devices by continuously monitoring previously-placed pages in HSS. Harmonia assigns a reward to the migration agent based on the impact of its migrations on future HSS performance. Second, we design Harmonia’s data-placement agent by extending the state-of-the-art data-placement technique, Sibyl (singh2022sibyl, ). To coordinate with the migration agent, we add new state features and improve reward structure for the data-placement agent. The two agents coordinate implicitly, i.e., they influence each other’s decisions through actions that affect the HSS environment. Unlike prior works (e.g., (rl_opti_TACO_2017, ; qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mao2023multi, ; mlsys2023qiu, ; lu2022rlrp, ; jain2017cooperative, ; jain2017coordinated, ; shen2022multiagent, )) that propose multi-agent RL approaches for various system optimizations and use (1) identical agents, and (2) a joint-action space and shared reward functions for all agents, Harmonia is the first multi-agent RL technique to co-optimize the different tasks of data placement and migration in HSS using heterogeneous agents with independent action spaces and reward structures.
@@
  • We demonstrate on a real Linux-based system with various HSS configurations that prior data-management approaches fall short of an Oracle policy due to the lack of combined optimization of data-placement and data-migration policies.
  • We propose Harmonia, a multi-agent RL-based data-management technique for HSS that optimizes data-placement and data-migration policies using two light-weight RL agents and achieves coordination between the agents to improve HSS performance.
  • We conduct rigorous evaluation of Harmonia on a real system with various HSS configurations of up to four devices, showing that Harmonia consistently outperforms state-of-the-art data-management techniques across a wide range of workloads.
@@ 2.1. Hybrid Storage Systems
Fig. 1 presents an overview of a hybrid storage system (HSS). An HSS has two key components: (1) multiple storage devices of varying characteristics (e.g., I/O latency, I/O throughput/bandwidth, and device capacity) and storage protocols (e.g., NVM Express (NVMe) (nvme, ) or SATA (sata, )), and (2) an HSS management layer that orchestrates data placement and data migration in the HSS.
@@
The HSS management layer has three key functions. First, it (1) provides a unified flat virtual address space for the applications to store data across all the HSS devices, (2) maps a virtual address to a logical block address (LBA) in one of the underlying storage devices, and (3) translates application I/O requests to the underlying storage devices using NVMe (nvme, ) or SATA (sata, ) protocols. The HSS management layer is similar to the md (mdlinux, ) feature in Linux, which provides software RAID functionality (RAID 0) when multiple storage devices are present. Second, the HSS management layer improves performance by (1) placing data (i.e., data placement) in the best-fit storage device for each incoming I/O request, and (2) rearranging data (i.e., data migration) across storage devices to handle changes in data access patterns after data placement and manage the faster devices effectively. Prior works (e.g., (singh2022sibyl, ; matsui2017design, ; ren2019archivist, )) demonstrate that data-placement and data-migration policies have a significant impact on HSS performance. Third, the HSS management layer stores the following metadata in the host DRAM: (1) virtual address to LBA mapping, (2) I/O request characteristics (e.g., request type, request size), and (3) storage device characteristics (e.g., device capacities and their current utilization). The HSS management layer is typically implemented in the kernel space of the host operating system.
@@ 2.3. Why RL for Data Management in HSS?
Prior work (singh2022sibyl, ) shows that RL outperforms heuristic- and supervised-learning-based approaches for data placement in HSS. RL is a good fit for HSS data management because of three key reasons.
@@
First, unlike supervised learning approaches (e.g.,  (doudali2019kleio, ; shetti2019machine, )), an RL approach can learn online without prior training using labeled data. Online learning helps the RL agent learn dynamic changes in workload access patterns and system conditions, and continuously adapt the data-placement and data-migration policies using system-level feedback. To demonstrate the dynamic changes in workload access patterns, we show the virtual addresses accessed and I/O request sizes during a 300-second execution of an example workload, ssd-02 (see §5 for evaluation methodology), in Figs. 3(a) and 3(b) respectively.
Figure 3. Virtual addresses accessed (top) and request sizes (bottom) captured during a 300-second execution of ssd-02 workload. Each virtual address accessed is represented as a green dot in the top figure.
We make two key observations: (1) the application accesses a broad range of virtual addresses ($1\times 10^{6}$ $1\times 10^{9}$) without fixed patterns. (2) the I/O request sizes vary from random accesses (i.e., 4KiB) to large sequential accesses (i.e., 128KiB). We observe such dynamic access pattern variations in all our chosen workloads. HSS system conditions can also change dynamically. For example, the device capacity utilization changes with each write operation, and access latencies vary significantly when multiple applications access a device concurrently. Prior heuristic- and supervised-learning-based approaches (see S3.1) make fixed placement decisions and perform migrations at infrequent intervals, and hence cannot adapt efficiently to the dynamic variations in workload access patterns and system conditions (see §3.2 for our motivational results).
@@
Second, RL can handle complex tasks such as data placement and data migration where each action changes the environment (e.g., device capacity utilization, bandwidth utilization, garbage collection) and impacts future outcomes (e.g., I/O request latencies). Unlike supervised learning approaches, RL can learn strategies that maximize long-term returns (e.g., (sutton_2018, ; sutton1999policy, ; ke2018modeling, )). Third, RL offers ease of extensibility to a wide range of HSS configurations and devices with heterogeneous characteristics with minimum designer effort.
We conclude that RL is a suitable approach to handle data management in HSS.
3. Motivation: Experimental Study on
Existing HSS Data-Management Techniques
3.1. Methodology
We evaluate eight prior data-management techniques (see Table 1) on a real Linux-based system with two HSS configurations, performance-optimized and cost-optimized, on 15 data-intensive workloads from three benchmark suites (see Table 5). We explain below the prior techniques evaluated, focusing on their key mechanisms. For more details on our evaluation methodology, see §5.
Table 1. Comparison of HSS Data-Management Techniques.
Technique
Data
Placement
Data
Migration
Key Details
CDE (matsui2017design, ) Heuristic Heuristic
Both policies are hand-tuned.
Data eviction during
I/O request handling.
RNN-HSS (doudali2019kleio, ) Heuristic
Supervised
Learning
Needs frequent retraining.
Data migration at fixed intervals.
K-SVM (shetti2019machine, ) Heuristic
Supervised
Learning
Needs frequent retraining.
Data migration at fixed intervals.
Sibyl (singh2022sibyl, ) RL Heuristic
Data eviction during
I/O request handling.
Sibyl (singh2022sibyl, )
+ K-SVM (shetti2019machine, )
RL
Supervised
Learning
Lack of coordination
between the two policies.
Sibyl (singh2022sibyl, )
+ RNN-HSS (doudali2019kleio, )
RL
Supervised
Learning
Lack of coordination
between the two policies.
CDE (matsui2017design, )
+ RL-Migr
Heuristic RL
Hand-tuned data-placement policy.
SAPM RL
Single RL agent for both tasks.
Harmonia RL RL
Two coordinating RL agents.

### v2

@@ Abstract.
Hybrid storage systems (HSS) combine multiple storage devices with diverse characteristics to achieve high performance and capacity at low cost. The performance of an HSS highly depends on the effectiveness of two key policies: (1) the data-placement policy, which determines the best-fit storage device for incoming data, and (2) the data-migration policy, which rearranges stored data (i.e., prefetches hot data and evicts cold data) across the devices to sustain high HSS performance. Prior works focus on improving only data placement or only data migration in HSS, which leads to relatively low HSS performance. Unfortunately, no prior work tries to optimize both policies together.
Our goal is to design a holistic data-management technique that optimizes both data-placement and data-migration policies to fully exploit the potential of an HSS, and thus significantly improve system performance. We demonstrate the need for multiple reinforcement learning (RL) agents to accomplish our goal. We propose Harmonia, a multi-agent reinforcement learning (RL)-based data-management technique that employs two lightweight autonomous RL agents, a data-placement agent and a data-migration agent, which adapt their policies for the current workload and HSS configuration, and coordinate with each other to improve overall HSS performance.
@@ 1. Introduction
A hybrid storage system (HSS) (xiao2016hs, ; wang2017larger, ; lee2014mining, ; felter2011reliability, ; bu2012optimization, ; krish2016efficient, ; lin2011hot, ; niu2018hybrid, ; tai2015sla, ; ou2014edm, ; cheng2015amc, ; zong2014faststor, ; matsui2017design, ; heuristics_hyrbid_hystor_sc_2011, ; lv2013probabilistic, ; guerra2011cost, ; elnably2012efficient, ; heuristics_usenix_2014, ; zhang2010automated, ; wu2012data, ; iliadis2015exaplan, ; lv2013hotness, ; matsui2017tri, ; feng2014hdstore, ; yang2017autotiering, ; kim2011hybridstore, ; liu2019hierarchical, ; luo2020optimal, ; singh2022sibyl, ; doudali2019kleio, ; ren2019archivist, ; cheng2019optimizing, ; shetti2019machine, ) is a cost-effective solution for high-performance and large-capacity storage requirements of data-intensive applications such as machine learning (DLRMFacebook, ; near_human_3, ; ho2020denoising, ), large language models (meta_llama, ; jiang2024mixtral, ; liu2024deepseek, ), databases (chan-signmod-1998, ; FastBitA9, ; redis-bitmaps, ), graph processing (beamer-SC-2012, ), and genome analysis (lander-nature-2001, ; altschul-jmb-1990, ; myers-jacm-1999, ). An HSS combines multiple storage devices (e.g., solid-state drives (SSDs) (inteloptane, ; intels4510, ; samsung2017znand, ; samsung-980pro, ; samsungmlc, ), hard disk drives (HDDs)) of diverse characteristics (e.g., I/O bandwidth, storage capacity). A typical HSS consists of fast storage devices with limited capacities and slow storage devices with large capacities.
@@
The performance of an HSS is highly dependent on two key policies, data placement and data migration. The data-placement policy determines the best-fit storage device in the HSS for incoming I/O requests. The data-migration policy rearranges data across the storage devices (i.e., prefetches frequently-accessed data to the fast device and evicts cold data to the slow device) to sustain high HSS performance, which can degrade over time due to (1) misplacement of data, and (2) changes in workload access patterns.
@@
We identify four key challenges in designing efficient data-place-ment and data-migration policies. First, workload access patterns and HSS conditions (e.g., access latencies, device capacity utilization) can change frequently in data-intensive environments, which makes it hard to optimize the policies. Second, the data-placement policy should have a low performance overhead as it operates on the critical path of I/O request handling. Third, the data-migration policy needs to migrate data across storage devices in a timely manner without impacting the latency of incoming I/O requests. Fourth, the two policies should not make conflicting decisions, which can adversely impact HSS performance and device lifetimes.
@@
Limitations of prior works. Prior works propose techniques for either data placement (e.g., (matsui2017design, ; sun2013high, ; heuristics_hyrbid_hystor_sc_2011, ; lv2013probabilistic, ; guerra2011cost, ; elnably2012efficient, ; heuristics_usenix_2014, ; bu2012optimization, ; krish2016efficient, ; tai2015sla, ; zhang2010automated, ; wu2012data, ; iliadis2015exaplan, ; lv2013hotness, ; matsui2017tri, ; feng2014hdstore, ; yang2017autotiering, ; singh2022sibyl, ; doudali2019kleio, ; ren2019archivist, ; cheng2019optimizing, ; shetti2019machine, ; li2017utility_HMM, ; agarwal2015page_HMM, ; agarwal2017thermostat_HMM, ; ham2013disintegrated_HMM, ; lin2011hot, ; salkhordeh2015operating, ; pavlovic2013data, ; meza2013case, ; chiachenchou2015batman, ; kim2011hybridstore, ; wang2019panthera, ; ramos2011page, ; liu2019hierarchical, ; luo2020optimal, ; doudali2021cori, ; sen2019machine, ; wu2019data, ; yang2014statistics, ; yolchuyev2021data, ; strunk2012hybrid, )) or data migration (e.g., (doudali2019kleio, ; lin2011hot, ; lin2017efficient, ; lin2018buffer, ; zhang2010adaptive, ; cheng2015amc, ; shetti2019machine, ; vengerov2008reinforcement, ; vasilakis2020hybrid2, )), but they provide relatively low performance when employed together or alone. Our motivational study on performance- (cost-) optimized HSS configurations demonstrates that even the state-of-the-art HSS data-placement technique, Sibyl (singh2022sibyl, ), only achieves 49.9% (41.4%) of the performance of an Oracle policy (see §3 for details). We identify two key reasons for such large performance gaps. First, no prior work proposes a holistic approach to optimize both data placement and data migration together. Second, heuristic- and supervised-learning-based policies (e.g., (singh2022sibyl, ; matsui2017design, ; matsui2017tri, ; doudali2019kleio, ; shetti2019machine, ; ren2019archivist, ; lin2011hot, ; lin2017efficient, )) require hand-tuning and may not adapt to changes in workload characteristics and HSS states.
@@
Since there are no prior works that optimize both placement and migration, we combine prior techniques to create four extended data-management techniques (see §3). Our evaluation shows that these extended techniques underperform because they are naive combinations of prior data-placement and data-migration approaches that are oblivious to each other’s decisions. We conclude that for high HSS performance, both data placement and migration policies need to coordinate with each other.
@@
Our goal is to design a holistic data-management technique for HSS that optimizes both data-placement and data-migration policies in a coordinated way to improve overall system performance. To this end, we propose Harmonia, [^1] Harmonia (wiki:harmonia, ): Greek goddess of harmony and balance, who brings peace to conflicts. a multi-agent online reinforcement learning (RL) based holistic approach to co-optimize HSS data placement and data migration.
@@
Need for multi-agent RL. We describe the benefits of RL for HSS data management in §3.1. Despite these benefits, a single RL agent cannot effectively (1) co-optimize (yang2020multi, ; yu2020meta, ; teh2017distral, ; sun2020adashare, ) different objectives like data placement and data migration, and (2) perform both tasks concurrently. Our motivational study demonstrates that a single-agent RL approach that performs both data placement and data migration leads to 18.8% (27.7%) lower average performance than Sibyl, which performs only data placement, on a performance- (cost-) optimized HSS (see §3.4). This is because learning to optimize data migration interferes with the process of learning an optimal data placement policy as they are two different tasks (See §3.2).
@@
Key Idea. Harmonia employs two coordinating lightweight autonomous RL agents, respectively for data placement and migration, to improve HSS performance. First, we design Harmonia’s data-placement agent by extending the state-of-the-art placement technique, Sibyl (singh2022sibyl, ). To coordinate with the migration agent, we add new state features and improve reward structure for the placement agent. Second, we design Harmonia’s data-migration agent to identify migration candidates and their target devices by continuously monitoring previously-placed pages in HSS. Harmonia assigns rewards to the migration agent based on the impact of its migrations on future HSS performance. The two agents influence each other’s decisions through actions that affect the HSS environment.
@@
This work makes the following key contributions:
  • We show that prior data-management approaches have low HSS performance because they do not optimize both data placement and data migration, and we quantify this on a real Linux system with various HSS configurations.
  • We propose Harmonia, a multi-agent RL-based data-management technique for HSS that co-optimizes data-placement and data-migration policies using two lightweight coordinating RL agents to improve HSS performance.
  • Via a rigorous real-system evaluation of Harmonia on various HSS configurations, we demonstrate that Harmonia consistently outperforms state-of-the-art data-management techniques across a wide range of data-intensive workloads.
@@ 2.1. Hybrid Storage Systems
Fig. 1 presents an overview of a hybrid storage system (HSS). An HSS has two key components: (1) multiple storage devices of varying characteristics (e.g., I/O latency, I/O bandwidth, and device capacity) and storage protocols (e.g., NVM Express (NVMe) (nvme, ) or SATA (sata, )), and (2) an HSS management layer that orchestrates data placement and data migration in the HSS.
@@
HSS Management Layer. The HSS management layer has three key functions. First, it (1) provides a unified flat virtual address space for the applications to store data across all HSS devices, (2) maps a virtual address to a logical block address (LBA) in one of the underlying storage devices, and (3) translates application I/O requests to the devices using NVMe (nvme, ) or SATA (sata, ) protocols. The HSS management layer is similar to the md (mdlinux, ) feature in Linux, which provides software RAID functionality (RAID 0) when multiple storage devices are present. Second, it (1) places data (i.e., data placement) in the best-fit storage device for each incoming I/O request, and (2) rearranges data (i.e., data migration) across storage devices. Prior works (e.g., (singh2022sibyl, ; matsui2017design, ; ren2019archivist, )) demonstrate that data-placement and migration policies have a significant impact on HSS performance. Third, the HSS management layer stores the following metadata in the host DRAM: (1) virtual address to LBA mapping, (2) I/O request characteristics (e.g., request type, request size), and (3) storage device characteristics (e.g., device capacities and their current utilization). The HSS management layer is typically implemented in the kernel space of the host OS.
@@ 2.2. Reinforcement Learning
2.3. Multi-Agent Reinforcement Learning
Multi-agent RL (e.g., (canese2021multi, ; rl_opti_TACO_2017, ; qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mao2023multi, ; mlsys2023qiu, ; lu2022rlrp, ; jain2017cooperative, ; jain2017coordinated, ; shen2022multiagent, )) is an extension of traditional RL where multiple RL agents interact with a shared environment and and with one another to learn optimal policies. Unlike RL with a single agent, multi-agent RL can effectively optimize multiple different tasks. A single RL agent cannot effectively learn optimal policies for different tasks because it relies on task similarity (ammar2014automated, ; carroll2005task, ) to transfer the learning gained from one task to another. Task similarity refers to the resemblance of tasks in their inputs, objectives, and actions. Prior works (yang2020multi, ; yu2020meta, ; teh2017distral, ; sun2020adashare, ; ammar2014automated, ; carroll2005task, ) show that a single RL agent cannot effectively learn optimal policies for multiple tasks with low task similarity because transferring the learning from one task to a dissimilar task can inhibit performance.
3. Motivation
In this section, we describe (1) why RL is a good fit for HSS data management, (2) the need for a multi-agent RL approach, and (3) prior approaches to HSS data management and their effectiveness.
@@ 3.1. Why RL for Data Management in HSS?
RL is a good fit for HSS data management for three key reasons.
@@
First, unlike heuristic and supervised-learning approaches, an RL approach can learn online without prior training using labeled data. Online learning helps the RL agent learn dynamic changes in workload access patterns and system conditions, and continuously adapt the data-placement and data-migration policies using system-level feedback without human intervention. Prior work Sibyl (singh2022sibyl, ) shows that RL outperforms heuristic-(e.g., (matsui2017design, ; matsui2017tri, )) and supervised-learning-based (e.g.,  (doudali2019kleio, ; shetti2019machine, )) approaches for HSS data placement. However, Sibyl partially uses RL for HSS data-management (i.e., RL for data-placement and a heuristic approach for data migration) that leaves a large gap between its performance and that of an Oracle approach (as we discuss in §3.4).
@@
Second, RL can handle complex tasks such as placement and migration where each action changes the environment (e.g., device capacity utilization, bandwidth utilization, garbage collection) and impacts future outcomes (e.g., I/O request latencies). Unlike supervised-learning approaches, RL learns strategies that maximize long-term returns (e.g., (sutton_2018, ; sutton1999policy, ; ke2018modeling, )), i.e., overall HSS performance.
Third, RL offers ease of extensibility to a wide range of HSS configurations and devices with heterogeneous characteristics with minimum designer effort (as partially demonstrated in Sibyl (singh2022sibyl, )).
3.2. Need for Multi-Agent RL for HSS
Data placement and data migration are two different tasks in HSS, and a multi-agent RL approach is a great fit for optimizing these different tasks (see §2.3).
Data placement and data migration in HSS are two different tasks for three key reasons. First, they operate on different data in the HSS. While data placement writes data from the current I/O request to the best-fit storage device, data migration monitors previously placed data to identify candidate pages for migration and their target storage devices. Second, their objectives are different. The objective of data placement is to minimize the I/O latency for each incoming request, but data migration aims to improve long-term HSS performance. Third, these two tasks are performed asynchronously with each other. Data placement is on the critical path of I/O handling and is performed synchronously with application writes or updates. Data migration is a background process and it is performed asynchronously to the application’s I/O requests.
To describe the need for data migration in addition to data placement, we provide two examples: (1) If a page in the fast device is rarely accessed, migrating it to the slow device can free up space for more frequently-accessed pages and improve overall HSS performance. (2) If a page on the slow device is accessed frequently (due to access pattern changes), it should be migrated to the fast device to reduce its access latency. In §3.4, we quantitatively demonstrate that a single RL agent (SAPM) shows limited performance when handling both data placement and migration in HSS. We conclude that multi-agent is a great fit for optimizing both data placement and migration in HSS, which are complex and different tasks.
3.3. Prior Approaches to HSS Data Management
Our focus in this work is to co-optimize placement and migration in HSS. In this section, we briefly introduce prior HSS data-management techniques. However, there is no existing technique that optimizes both placement and migration in HSS. We discuss different approaches to optimize both placement and migration by combining (or extending) prior HSS data-management techniques.